# Identifying Compton-thick AGNs in the COSMOS

### v1

@@ 4.1 Multiwavelength photometric data
The multiwavelength photometric data used in this work span the far-UV (FUV, 1526 Å) to the far-infrared (FIR, 500 m). Among them, the photometric data covering the wavelength range from the FUV to the MIR (27 bands in total) are sourced from the COSMOS2020 FARMER catalog (Weaver et al., 2022). These bands include the GALEX FUV and NUV band, MegaCam U band, CFHT U band, five Subaru HSC bands (g, r, i, z, and y), 12 Subaru Suprime-Cam medium bands (IB427, IB464, IA484, IB505, IA527, IB574, IA624, IA679, IB709, IA738, IA767, and IB827), four VISTA VIRCAM broad bands (Y, J, H, and ), and four Spitzer IRAC bands (ch1, ch2, ch3, and ch4). Moreover, the FIR photometric data are sourced from the COSMOS2015 catalog (Laigle et al., 2016), including MIPS 24 m, PACS 100 m, PACS 160 m, SPIRE 250 m, SPIRE 350 m, and SPIRE 500 m. Where photometric data in a given band are missing, we use the corresponding flux upper limits in the SED analysis.
@@ 4.2 Analysis of SEDs
The SED analysis of AGNs is a widely used technique that not only provides detailed information about AGNs but also reveals the properties of their host galaxies (e.g., Guo et al., 2020), such as AGN luminosity, rest-frame 6 m luminosity for the AGN, stellar mass (M⋆), and the star formation rate (SFR). In this work, we use Code Investigating GALaxy Emission (CIGALE; V2022.1; Boquien et al., 2019), an open Python code containing the template of galaxies and AGNs, to conduct the analysis of AGNs in our sample and obtain the required parameters.
Since all sources in our sample are MIR-selected AGNs, we have used templates that incorporate both the galaxy and AGN characteristics to achieve the most accurate SED fitting. The galaxy templates in CIGALE are generated by integrating four distinct modules, including the star formation history (SFH), the single stellar population (SSP) model (Bruzual & Charlot, 2003), the dust attenuation (DA; Calzetti et al., 2000), and the dust emission (DE; Draine & Li, 2007; Dale et al., 2014). The AGN component in our analysis is a module developed by Fritz et al. (2006). The modules and parameters used for SED fitting are summarized in Table 1. Additionally, for sources using upper flux limits in certain bands, the lim_flag is set to ”noscaling”. Given the vast parameter space, quickly constructing a model on the computer is not feasible. Thus, we adopted an iterative methodology to expedite the acquisition of the best-fitting SEDs. Fig. 3 shows the best-fitting SED for an AGN. To assess the reliability of the fitted AGN components, we refit SEDs using only the galaxy templates. A markedly better fit (indicated by a lower ) with the combined galaxy+AGN templates than with the galaxy-only templates implies that the fitted AGN components can indeed represent the contribution of the actual AGNs in their SEDs. We used the Bayesian information criterion (BIC) to quantify the significance or confidence of the AGN components (see Appendix C of Guo et al., 2025, for details).
@@
Through SED fitting, we obtained the best-fitting SED for each source and derived the corresponding physical parameters. Cols. 9–10 of Table 2 are rest-frame 6 m luminosity for the AGN and the AGN components’ confidence. Col. 11 of Table 2 presents M⋆, the derived outcome of SED fitting. However, since the SFR derived from SED fitting relies on the chosen SFH model (Gao et al., 2019), we instead use the calibration method proposed by Bell et al. (2005) to estimate the SFR using UV and IR luminosities. This calibration is scaled to the Chabrier (2003) initial mass function for accuracy. Specifically, the SFR is calculated using the formula
where is an estimation of the integrated 1216 – 3000 Å rest-frame UV luminosity, and is the 8 – 1000 m rest-frame IR luminosity. Both and are in units of . Col. 12 of Table 2 lists the SFR of each source.
@@ 5.1 MIR diagnostics
The emissions from the corona and dust torus of AGNs are considered good tracers of the accretion disk’s emission. Consequently, a strong correlation is expected between X-ray and MIR emissions in AGNs. Several studies have investigated this relationship in radio-quiet AGNs (e.g., Fiore et al., 2009; Mateos et al., 2015; Stern, 2015; Chen et al., 2017; Esparza-Arredondo et al., 2019, 2021). Due to the low optical depth in the MIR band, the MIR radiation from CT-AGNs is not significantly suppressed. As the column density increases, the torus absorbs more nuclear X-ray emission and re-radiates it in the MIR, potentially boosting the MIR luminosity. However, this effect is extremely weak: the AGN X-ray luminosity is 1–2 dex below its UV–optical continuum (Steffen et al., 2006; Lusso et al., 2010), so even complete absorption and re-radiation in the MIR would contribute negligibly compared with the direct disk-dust heating. Ricci et al. (2017b) indeed found no significant increase in MIR luminosity with column density. Consequently, the intrinsic X-ray and MIR luminosities of AGNs are still expected to follow the above relations. However, because the X-ray radiation of CT-AGNs is heavily absorbed, their observed X-ray luminosities fall below the values predicted by these relations. Given that the AGNs in our sample are not detected in X-rays, we adopt the corresponding X-ray luminosity upper limits as their observed luminosities and use these relationships to identify CT-AGNs.
@@
Fig. 4 shows the observed X-ray luminosity in the rest-frame 2–10 keV band as a function of the 6 m luminosity for the fit AGN component. The solid purple, blue, and black lines represent the relation for Chen et al. (2017), Stern (2015), and Fiore et al. (2009), respectively. Assuming the intrinsic 2–10 keV spectrum of each AGN is modeled in in Xspec with phabspowerlaw+constantpowerlaw+pexmon (see Appendix D of Guo et al. 2025 for details), the dashed lines represent the same relationships between MIR and X-ray luminosities after the X-ray spectra have been absorbed by gas with a column density of . This means that the sources within or below the shaded area in Fig. 4 should be classified as CT-AGNs. Circles tipped with downward arrows denote the AGNs in our sample, their colors indicating the confidence levels of the fitted AGN components. As shown in Fig. 4, 23 AGNs fall within or below the shaded area, suggesting that they may be CT-AGNs. However, considering the variability in the relations provided by Chen et al. (2017) and Fiore et al. (2009), a range of possible CT-AGN candidates is 7 to 23. To further determine whether they are CT-AGNs, we must assess the reliability of their fitted AGN components. These sources, whose fitted AGN components exhibit a high confidence coefficient (), support a CT-AGN classification, whereas those with a low confidence coefficient () cannot be reliably identified as CT-AGNs. Finally, a total of 23 AGNs located within or below the shaded region exhibit high confidence coefficients, supporting their diagnosis as CT-AGNs. These CT-AGNs are marked in Fig. 4 by cutting stars within the large circles. Col. 13 of Table 2 also lists their classification, ”1” represents CT-AGN identified in this work.
@@ 5.2 Stacking Analysis for CT-AGNs
In the preceding section, we successfully find out 7 to 23 CT-AGN candidates. Because these AGNs are individually undetected in X-rays, we cannot extract their X-ray spectra and determine their column densities. To confirm heavy absorption in the X-ray band, we perform X-ray stacking analysis on these 23 AGNs using the publicly available CSTACK V4.5 tool (Miyaji et al., 2008)333CSTACK was developed by Takamitsu Miyaji and is available at https://lambic.astrosen.unam.mx/cstack/. This analysis adopts the default maximum off-axis angle, source aperture, background annulus, and automatic exclusion of known X-ray sources. The output of CSTACK includes the stacked photon count rates and their errors for both the soft (0.5–2 keV) and hard (2–8 keV) bands, along with the corresponding stacked images. Fig. 5 presents the stacking results: the stacked AGNs are clearly detected at significance in the soft band and marginally detected at in the hard band. The stacked photon count rates in the soft and hard bands are and , respectively.
@@
To further constrain the absorption of these sources, we fitted the soft- and hard-band fluxes in Xspec with the zphabszpowerlw+constantzpowerlw+pexmon model. The redshift was fixed at the mean value of the 23 sources (z = 0.92), the photon index at 1.9, the inclination angle at 85, and the scattered fraction at 0.005; only the column density and normalisation norm were allowed to vary, while all remaining parameters were kept at their default values. Fig. 6 presents the p-value distribution in the – norm plane obtained by fitting the soft- and hard-band fluxes with our model. In – norm plane, the data are consistent with the model at the 95% confidence level within two distinct parameter regions, corresponding to Compton-thin and Compton-thick solutions, respectively. The golden solid circles mark the best-fitting points in both regions, with corresponding column densities of and , respectively. The X-ray upper limits of these 23 sources already fall below the relationships between MIR and X-ray luminosities after absorbing by Compton-thick material. If their absorption were located in the Compton-thin region, the majority of them would have to be intrinsically X-ray weak AGNs. However, the intrinsic incidence of X-ray weak AGNs is far lower than that of CT-AGNs, making it extremely improbable that most of them are X-ray weak AGNs. The Compton-thin solution can therefore be ruled out. Within Compton-thick region, the 95% confidence level demands , while the 90% confidence level requires . Both thresholds lie well above the Compton-thick limit, providing direct and compelling evidence that these sources are CT-AGNs.
@@ 5.3 Missed Diagnosis of CT-AGNs
Although MIR-selected AGN samples are incomplete, numerous studies have shown that their CT-AGN fraction closely matches—or even reaches—the theoretical expectation of CXB (e.g., Lyu et al., 2024; Boorman et al., 2025; Annuar et al., 2025). Our MIR-selected sample excludes X-ray-detected AGNs, which preferentially select less absorbed AGNs. Consequently, we expect the CT-AGN fraction in our MIR-selected sample to markedly exceed that predicted by CXB models. However, our MIR diagnostics identified only 23 CT-AGNs, 2.1% (23/1104) of the MIR-selected sample, significantly below the fraction expected from CXB synthesis models. This discrepancy clearly exceeds statistical error, indicating that MIR diagnostics miss a large number of CT-AGNs. Based on our analysis of the CT-AGN fraction within our sample, we infer that at least 300 CT-AGNs remain unidentified. With the current data and methods, we are still unable to find these missing CT-AGNs. To identify these missed CT-AGNs, the most direct and efficient approach is to conduct significantly deeper X-ray exposures of the COSMOS in the future. On the one hand, deeper exposures will enhance detection sensitivity, revealing AGNs that are currently buried in the background. This may not only confirm their existence but also allow column density measurements via X-ray spectral fitting, thereby reliably identifying CT-AGNs. On the other hand, for AGNs that remain undetected even after deep exposures, longer integrations will push their X-ray luminosity upper limits to lows, markedly enhancing the discriminatory power of MIR diagnostics.
@@ 5.4 The properties of CT-AGN host galaxies
CT-AGNs are widely regarded as an early, heavily obscured phase of AGN evolution (e.g., Guo et al., 2023). Their growth is intimately coupled to the host-galaxy environment, as evidenced by the tight MBH– relation (e.g., Ferrarese & Merritt, 2000) and the role of AGN feedback in regulating star formation (e.g., Bower et al., 2006), which together support a coevolutionary scenario between AGNs and their host galaxies (e.g., Kormendy & Ho, 2013). CT-AGNs are considered to preferentially host in galaxies with a significant amount of dense gas (Kocevski et al., 2015; Ricci et al., 2017a). Intense AGN radiation compresses their surrounding dense gas, potentially triggering star formation activity within their host galaxies. The host galaxies of CT-AGNs in the local universe exhibit a high level of star formation (e.g., Hopkins et al., 2006; Goulding et al., 2012). Zhang et al. (2025) find that the host galaxies of CT-AGNs in the CDFS exhibit higher levels of star formation activity than those of non-CT-AGNs. Andonie et al. (2022) suggest that obscured AGNs have star formation rates times higher than unobscured systems in the COSMOS. However, Gilli et al. (2022) suggest that the host galaxies of high-redshift AGNs are different from those of AGNs in the local universe. Guo et al. (2025) report similar stellar mass and SFR distributions for CT-AGN and non-CT-AGN host galaxies; while their sample contains only 18 CT-AGNs, which is insufficient for statistically robust conclusions. Using a substantially larger sample, we will re-examine the SFRs and stellar masses of CT-AGN host galaxies relative to those of non-CT-AGNs.
Because a significant number of CT-AGNs in our sample remain unidentified, we cannot extract a pure non-CT-AGN sample. We therefore adopt the non-CT-AGNs provided by Guo et al. (2025) as the control sample, and combine their CT-AGNs with the 23 newly identified in this work to construct the final CT-AGN sample. Fig. 7 presents relationships, known as the main sequence, between SFRs and stellar masses of the host galaxies for both CT-AGNs and non-CT-AGNs. A Kolmogorov–Smirnov test (KS test) of stellar mass distributions for their host galaxies gives ; repeating the KS test on their SFR distributions yields . Although these values are slightly lower than those reported by Guo et al. (2025), they still indicate no significant difference between the two distributions. Our findings, therefore, do not support a possible correlation between intense star formation and CT-AGNs reported by Zhang et al. (2025). This finding is likely due to the limited size of our CT-AGN sample, which is still too small for statistically robust conclusions. In future work, we will enlarge the CT-AGN sample and further study whether thereare differences in the host galaxies of both CT-AGNs and non-CT-AGNs.

### v2

@@ 4.1 Multiwavelength photometric data
The multiwavelength photometric data used in this work span the far-UV (FUV; 1526 Å) to the far-infrared (FIR; 500 m). Among them, the photometric data covering the wavelength range from the FUV to the MIR (27 bands in total) are sourced from the COSMOS2020 FARMER catalog (Weaver et al., 2022). These bands include the GALEX FUV and NUV band, MegaCam U band, CFHT U band, five Subaru HSC bands (g, r, i, z, and y), 12 Subaru Suprime-Cam medium bands (IB427, IB464, IA484, IB505, IA527, IB574, IA624, IA679, IB709, IA738, IA767, and IB827), four VISTA VIRCAM broad bands (Y, J, H, and ), and four Spitzer IRAC bands (ch1, ch2, ch3, and ch4). Moreover, the FIR photometric data are sourced from the COSMOS2015 catalog (Laigle et al., 2016), including MIPS 24 m, PACS 100 m, PACS 160 m, SPIRE 250 m, SPIRE 350 m, and SPIRE 500 m. Where photometric data in a given band are missing, we use the corresponding flux upper limits in the SED analysis.
@@ 4.2 Analysis of SEDs
The SED analysis of AGNs is a widely used technique that not only provides detailed information about AGNs but also reveals the properties of their host galaxies (e.g., Guo et al., 2020), such as AGN luminosity, rest-frame 6 m luminosity for the AGN, stellar mass (M⋆), and the star formation rate (SFR). To conduct the analysis of AGNs in our sample and obtain the required parameters in this work, we used Code Investigating GALaxy Emission (CIGALE; V2022.1; Boquien et al., 2019), an open Python code containing the template of galaxies and AGNs. Since all the sources in our sample are MIR-selected AGNs, we have used templates that incorporate both the galaxy and AGN characteristics to achieve the most accurate SED fitting. The galaxy templates in CIGALE are generated by integrating four distinct modules, including the star formation history (SFH), the single stellar population (SSP) model (Bruzual & Charlot, 2003), the dust attenuation (DA; Calzetti et al., 2000), and the dust emission (DE; Draine & Li, 2007; Dale et al., 2014). The AGN component in our analysis is a module developed by Fritz et al. (2006). The modules and parameters used for SED fitting are summarized in Table 1. Additionally, for sources using upper flux limits in certain bands, the lim_flag is set to ”noscaling.” Given the vast parameter space, quickly constructing a model on the computer is not feasible. Thus, we adopted an iterative methodology to expedite the acquisition of the best-fitting SEDs. Figure 3 shows the best-fitting SED for an AGN. To assess the reliability of the fitted AGN components, we refit SEDs using only the galaxy templates. A markedly better fit (indicated by a lower ) with the combined galaxy+AGN templates than with the galaxy-only templates implies that the fitted AGN components can indeed represent the contribution of the actual AGNs in their SEDs. We used the Bayesian information criterion (BIC) to quantify the significance or confidence of the AGN components (see Appendix C of Guo et al., 2025, for details).
@@
Through SED fitting, we obtained the best-fitting SED for each source and derived the corresponding physical parameters. Columns 9–10 of Table 2 are rest-frame 6 m luminosity for the AGN and the AGN components’ confidence. Column 11 of Table 2 presents M⋆, the derived outcome of SED fitting. However, since the SFR derived from SED fitting relies on the chosen SFH model (Gao et al., 2019), we instead use the calibration method proposed by Bell et al. (2005) to estimate the SFR using UV and IR luminosities. This calibration is scaled to the Chabrier (2003) initial mass function for accuracy. Specifically, the SFR is calculated using the formula,
where is an estimation of the integrated 1216 – 3000 Å rest-frame UV luminosity, and is the 8 – 1000 m rest-frame IR luminosity. Both and are in units of . Column 12 of Table 2 lists the SFR of each source.
@@ 5.1 MIR diagnostics
The emissions from the corona and dust torus of AGNs are considered good tracers of the accretion disk’s emission. Consequently, a strong correlation is expected between X-ray and MIR emissions in AGNs. Several studies have investigated this relationship in radio-quiet AGNs (e.g., Fiore et al., 2009; Mateos et al., 2015; Stern, 2015; Chen et al., 2017; Esparza-Arredondo et al., 2019, 2021). Due to the low optical depth in the MIR band, the MIR radiation from CT-AGNs is not significantly suppressed. As the column density increases, the torus absorbs more nuclear X-ray emission and re-radiates it in the MIR, potentially boosting the MIR luminosity. However, this effect is extremely weak: the AGN X-ray luminosity is 1–2 dex below its UV–optical continuum (Steffen et al., 2006; Lusso et al., 2010), so even complete absorption and re-radiation in the MIR would contribute negligibly compared with the direct disk-dust heating. Ricci et al. (2017b) indeed found no significant increase in MIR luminosity with column density. Consequently, the intrinsic X-ray and MIR luminosities of AGNs are still expected to follow the above relations. However, because the X-ray radiation of CT-AGNs is heavily absorbed, their observed X-ray luminosities fall below the values predicted by these relations. Given that the AGNs in our sample are not detected in X-rays, we adopted the corresponding X-ray luminosity upper limits as their observed luminosities and use these relationships to identify CT-AGNs.
@@
Figure 4 shows the observed X-ray luminosity in the rest-frame 2–10 keV band as a function of the 6 m luminosity for the fit AGN component. The solid purple, blue, and black lines represent the relation for Chen et al. (2017), Stern (2015), and Fiore et al. (2009), respectively. Assuming the intrinsic 2–10 keV spectrum of each AGN is modeled in in Xspec with phabspowerlaw+constantpowerlaw+pexmon (see Appendix D of Guo et al. 2025 for details), the dashed lines represent the same relationships between MIR and X-ray luminosities after the X-ray spectra have been absorbed by gas with a column density of . This means that the sources within or below the shaded area in Fig. 4 should be classified as CT-AGNs. Circles tipped with downward arrows denote the AGNs in our sample, their colors indicating the confidence levels of the fitted AGN components. As shown in Fig. 4, 23 AGNs fall within or below the shaded area, suggesting that they may be CT-AGNs. However, considering the variability in the relations provided by Chen et al. (2017) and Fiore et al. (2009), a range of possible CT-AGN candidates is 7 to 23. To further determine whether they are CT-AGNs, we must assess the reliability of their fitted AGN components. These sources, whose fitted AGN components exhibit a high confidence coefficient (), support a CT-AGN classification; whereas those with a low confidence coefficient () cannot be reliably identified as CT-AGNs. Finally, a total of 23 AGNs located within or below the shaded region exhibit high confidence coefficients, supporting their diagnosis as CT-AGNs. These CT-AGNs are marked in Fig. 4 by cutting stars within the large circles. Column 13 of Table 2 also lists their classification, ”1” represents CT-AGN identified in this work.
@@ 5.2 Stacking analysis for CT-AGNs
In the preceding section, we successfully find out 7 to 23 CT-AGN candidates. Because these AGNs are individually undetected in X-rays, we cannot extract their X-ray spectra and determine their column densities. To confirm heavy absorption in the X-ray band, we perform X-ray stacking analysis on these 23 AGNs using the publicly available CSTACK V4.5 tool (Miyaji et al., 2008)333CSTACK was developed by Takamitsu Miyaji and is available at https://lambic.astrosen.unam.mx/cstack/. This analysis adopts the default maximum off-axis angle, source aperture, background annulus, and automatic exclusion of known X-ray sources. The output of CSTACK includes the stacked photon count rates and their errors for both the soft (0.5–2 keV) and hard (2–8 keV) bands, along with the corresponding stacked images. Figure 5 presents the stacking results: the stacked AGNs are clearly detected at significance in the soft band and marginally detected at in the hard band. The stacked photon count rates in the soft and hard bands are and , respectively.
@@
To further constrain the absorption of these sources, we fitted the soft- and hard-band fluxes in Xspec with the zphabszpowerlw+constantzpowerlw+pexmon model. The redshift was fixed at the mean value of the 23 sources (z = 0.92), the photon index at 1.9, the inclination angle at 85, and the scattered fraction at 0.005; only the column density, , and normalization norm were allowed to vary, while all remaining parameters were kept at their default values. Figure 6 presents the p-value distribution in the – norm plane obtained by fitting the soft- and hard-band fluxes with our model. In – norm plane, the data are consistent with the model at the 95% confidence level within two distinct parameter regions, corresponding to Compton-thin and Compton-thick solutions, respectively. The golden solid circles mark the best-fitting points in both regions, with corresponding column densities of and , respectively. The X-ray upper limits of these 23 sources already fall below the relationships between MIR and X-ray luminosities after absorbing by Compton-thick material. If their absorption were located in the Compton-thin region, the majority of them would have to be intrinsically X-ray weak AGNs. However, the intrinsic incidence of X-ray weak AGNs is far lower than that of CT-AGNs, making it extremely improbable that most of them are X-ray weak AGNs. The Compton-thin solution can therefore be ruled out. Within Compton-thick region, the 95% confidence level demands , while the 90% confidence level requires . Both thresholds lie well above the Compton-thick limit, providing direct and compelling evidence that these sources are CT-AGNs. As the hard X-ray band is only marginally detected (slightly above ) in the stacked spectrum, we conservatively treated its flux as an upper limit. Fitting the same model under this assumption still allows us to obtain a solution consistent with Compton-thick absorption.
@@ 5.3 Missed diagnosis of CT-AGNs
Although MIR-selected AGN samples are incomplete, numerous studies have shown that their CT-AGN fraction closely matches (and even has the potential to reach) the theoretical expectation of CXB (e.g., Lyu et al., 2024; Boorman et al., 2025; Annuar et al., 2025). Our MIR-selected sample excludes X-ray-detected AGNs, which preferentially select less absorbed AGNs. Consequently, we expect the CT-AGN fraction in our MIR-selected sample to markedly exceed that predicted by CXB models. However, our MIR diagnostics identified only 23 CT-AGNs, 2.1% (23/1104) of the MIR-selected sample, significantly below the fraction expected from CXB synthesis models. This discrepancy clearly exceeds statistical error, indicating that MIR diagnostics miss a large number of CT-AGNs. Based on our analysis of the CT-AGN fraction within our sample, we infer that at least 300 CT-AGNs remain unidentified. With the current data and methods, we are still unable to find these missing CT-AGNs. To identify these missed CT-AGNs, the most direct and efficient approach is to conduct significantly deeper X-ray exposures of the COSMOS in the future. On the one hand, deeper exposures will enhance detection sensitivity, revealing AGNs that are currently buried in the background. This might not only confirm their existence, but also enable column density measurements via X-ray spectral fitting, thereby reliably identifying CT-AGNs. On the other hand, for AGNs that remain undetected even after deep exposures, longer integrations will push their X-ray luminosity upper limits to lows, markedly enhancing the discriminatory power of MIR diagnostics.
@@ 5.4 The properties of CT-AGN host galaxies
CT-AGNs are widely regarded as an early, heavily obscured phase of AGN evolution (e.g., Guo et al., 2023). Their growth is intimately coupled to the host-galaxy environment, as evidenced by the tight MBH– relation (e.g., Ferrarese & Merritt, 2000) and the role of AGN feedback in regulating star formation (e.g., Bower et al., 2006). Together, this picture supports a coevolutionary relationship between AGNs and their host galaxies (e.g., Kormendy & Ho, 2013). CT-AGNs are considered to preferentially host in galaxies with a significant amount of dense gas (Kocevski et al., 2015; Ricci et al., 2017a). Intense AGN radiation compresses their surrounding dense gas, potentially triggering star formation activity within their host galaxies. The host galaxies of CT-AGNs in the Local Universe exhibit a high level of star formation (e.g., Hopkins et al., 2006; Goulding et al., 2012). Zhang et al. (2025) find that the host galaxies of CT-AGNs in the CDFS exhibit higher levels of star formation activity than those of non-CT-AGNs. Andonie et al. (2022) suggested that obscured AGNs have star formation rates that are about three times higher than unobscured systems in the COSMOS. However, Gilli et al. (2022) suggested that the host galaxies of high-redshift AGNs are different from those of AGNs in the Local Universe. Guo et al. (2025) reported similar stellar mass and SFR distributions for CT-AGN and non-CT-AGN host galaxies; while their sample contains only 18 CT-AGNs, which is insufficient for statistically robust conclusions. Using a substantially larger sample, we will re-examine the SFRs and stellar masses of CT-AGN host galaxies relative to those of non-CT-AGNs.
Because a significant number of CT-AGNs in our sample remain unidentified, we were not able to extract a pure non-CT-AGN sample. We therefore adopted the non-CT-AGNs provided by Guo et al. (2025) as the control sample and then combined their CT-AGNs with the 23 newly identified in this work to construct the final CT-AGN sample. Figure 7 presents relationships, represented by the main sequence, between SFRs and stellar masses of the host galaxies for both CT-AGNs and non-CT-AGNs. A Kolmogorov–Smirnov test (KS test) of stellar mass distributions for their host galaxies gave us , while repeating the KS test on their SFR distributions yielded . Although these values are slightly lower than those reported by Guo et al. (2025), they still indicate that there is no significant difference between the two distributions. Thus, our findings do not support any correlation between intense star formation and CT-AGNs, as reported by Zhang et al. (2025). This finding is likely due to the limited size of our CT-AGN sample, which is still too small for statistically robust conclusions. In future works, we will enlarge the CT-AGN sample and further study whether there are differences in the host galaxies of both CT-AGNs and non-CT-AGNs.